Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 236     Entry time: Tue Mar 5 14:42:22 2019
Author: Drive Team- Ino&Thierry 
Type: Telescope Operations 
Category: General 
Subject: High speed drive tests 
Status:  

. Fisrt day of the high speed drive tests with IFAE team.

  • Tests for tuning and generating motor plots at 10% (0.0234 Rad/s) of the maximal speed (GRB alert speed)
  • Tuning of the emergency stop from 0.05 Rad/s2 up to 0.2 Rad/s2
  • 6 movements at 25% speed. (from -90 to 270)
  • repositionnning to -90 at 40% 
  • movements from -90 to 270 and reverse at 40%
  • 3 movements from -90 to 270 and reverse at 60%  : the FW generates an alarm and the drive stops the movement.
    • The FW level was ok but the the status (GRB Fast Rotation and UPS system) got red.
    • Koji has been contacted to look at this issue.
ELOG V3.1.5-fc6679b